Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by bburton11 » Fri Jun 26, 2009 11:42 am
There are 15 slate rocks, 20 pumie rocks, and 10 granite rocks randomly distributed in a certain field. If 2 rocks are to be chosen at random and with replacement, what is the probability that both rocks will be slate rocks?

1/3
7/22
1/9
7/66
2/45

Am I missing something here. I tired all the ways i know to solve this probability problem however, I am need help.
Join the discussion
Source: — Problem Solving |

by DanaJ » Fri Jun 26, 2009 12:03 pm
To my knowledge, "with replacement" means that rock selected for your first draw is put back in the lot. This makes the second draw independent from the first, as explained here:

https://math.youngzones.org/joint.html

The probability that the first draw is a slate rock is 15/45 (number of favorable cases/number of possible cases) or 1/3. The probability of the second rock being a slate rock will be 1/3 as well, since we have replaced the initial draw in the lot (i.e. we've put it back in the set).
The probability of both events taking place is 1/3 * 1/3 or 1/9.

HOWEVER

If you consider a "without replacement" probability, then you get something different. The first draw will have the same probability, i.e. 1/3. The second draw will not, since the number of favorable cases has dropped from 15 to 14 (since you've drawn a slate rock in your first pick) and the total number of cases has dropped from 45 to 44. So for your second draw, you get probability of = 14/44 = 7/22.
Multiply the two to get 7/66.

What is the source of this problem? Not all sources are 100% error-proof. Neither is my explanation, so I'm waiting for other opinions as well.
Join the discussion

Re: Probability

by ssmiles08 » Fri Jun 26, 2009 12:19 pm
bburton11 wrote:There are 15 slate rocks, 20 pumie rocks, and 10 granite rocks randomly distributed in a certain field. If 2 rocks are to be chosen at random and with replacement, what is the probability that both rocks will be slate rocks?

1/3
7/22
1/9
7/66
2/45

Am I missing something here. I tired all the ways i know to solve this probability problem however, I am need help.
Hey bburton11,

are you sure the question says with replacement? or without replacement. I agree with DanaJ, b/c with replacement it would be (1/3)*(1/3) = 1/9

but w/o replacement... it would be 1/3*14/44 = 7/66.

also..If it does say with replacement, I am a bit suspicious b/c it doesn't say it is replaced with the same type of stone...it could have been replaced with a granite rock for all we know.
Join the discussion

by rahulg83 » Sat Jun 27, 2009 1:14 am
I agree, should be 1/9
Join the discussion